# On von Neumann regularity of ample groupoid algebras

### v2

@@ Abstract.
We study the question of when the algebra of an ample groupoid over a field is von Neumann regular and, more generally, when the algebra of a graded ample groupoid is graded von Neumann regular. For fields of characteristic 00 we are able to provide a complete characterization. Over fields of positive characteristic, we merely have some sufficient and some necessary conditions for regularity. As applications we recover known results on regularity and graded regularity of Leavitt path algebras and inverse semigroups. We also prove a number of new results, in particular concerning graded regularity of algebras of Deaconu-Renault groupoids and Nekrashevych-Exel-Pardo algebras of self-similar groups.
@@ 1. Introduction
A famous result of Connell [12] says that a group ring $KG$, over a field $K$, is regular if and only if $G$ is locally finite and the characteristic of $K$ does not divide the order of any element of $G$. In particular, this says that regularity is a rare feature for groups rings. Leavitt path algebras [1] are also rarely regular: they are regular if and only if the (directed) graph defining the Leavitt path algebra is acyclic [3]. However, every Leavitt path algebra is graded von Neumann regular by a result of Hazrat [20]. A graded ring $R$ is graded regular if the semigroup of homogeneous elements is regular. This has an important impact on the graded $K$-theory of Leavitt path algebras.
Okniński tried to generalize Connell’s results to inverse semigroup algebras (see his book [28]). He was able to prove that if $K$ is a field of characteristic 00 and $S$ is an inverse semigroup, then $KS$ is regular if and only if $S$ is locally finite. In positive characteristic he was able to give necessary conditions: $S$ is periodic, each maximal subgroup of $S$ is locally finite and the characteristic of $K$ cannot divide the order of any element of a maximal subgroup. It had been proven earlier that if $S$ is locally finite and the characteristic of $K$ divides the order of no element of any maximal subgroup, then $KS$ is regular. But being periodic with locally finite maximal subgroups is a strictly weaker condition than being locally finite. The reason that Okniński was able to obtain a complete characterization in characteristic 00 is that he could embed $\mathbb{C}S$ into the Banach algebra $\ell_{1}(S)$ and use some analysis.
@@
Our main result is a complete characterization of regularity of $K\mathscr{G}$ when $K$ is a field of characteristic 00. Following the Bourbaki tradition, we say a space is quasi-compact if every open covering has a finite subcover and reserve “compact” to mean quasi-compact and Hausdorff. Then $K\mathscr{G}$ is regular if and only if $\mathscr{G}$ is a directed union of quasi-compact open subgroupoids. For principal, second countable Hausdorff groupoids this is equivalent to approximate finiteness in the sense of [18], hence we call such groupoids approximately quasi-compact. Like Okniński, we use analytic methods, but this time we embed $\mathbb{C}\mathscr{G}$ into the reduced $C^{*}$-algebra $C^{*}_{r}(\mathscr{G})$.
In positive characteristic, we are able to show that if $\mathscr{G}$ is approximately quasi-compact and the characteristic of $K$ does not divide the order of any isotropy element of $\mathscr{G}$, then $K\mathscr{G}$ is regular. But we are unable to prove in general the necessity of being approximately quasi-compact in positive characteristic.
We also consider graded regularity of groupoid algebras for ample groupoids equipped with a locally constant cocycle to a group. We prove the algebra is graded regular if and only if the homogeneous component of the identity is regular. Since the homogeneous component of the identity is the algebra of the groupoid of elements mapping to the identity under the cocycle, our results for regularity of groupoid algebras apply. From this we easily recover Hazrat’s result on graded regularity of Leavitt path algebras [20]. We extend these results to algebras of Deaconu-Renault groupoids under mild hypotheses. These include algebras of row finite higher rank graphs. We consider further graded regularity of partial group skew products, Leavitt path algebras of labeled graphs and Nekrashevych-Exel-Pardo algebras of self-similar graphs.
The paper is organized as follows. Section 2 is devoted to preliminaries concerning inverse semigroups, ample groupoids and regular rings. Then Section 3 proves our main results on regularity of ample groupoid algebras. Section 4 addresses the question of graded regularity of the algebra of an ample groupoid equipped with a locally constant cocycle. The final section, applies the main results to study regularity and graded regularity for the algebras of a number of well-studied families of ample groupoids.
@@ 2.1. Inverse semigroups and groupoids
The reader is referred to [22] for details on the theory of inverse semigroups. A semigroup is a set together with an associative binary operation. An inverse semigroup is a semigroup $S$ such that for every $s\in S$ there is a unique $s^{*}\in S$ such that $s^{*}ss^{*}=s^{*}$ and $ss^{*}s=s$. Notice that $ss^{*},s^{*}s$ are idempotents for all $s\in S$. The set $E(S)$ of idempotents of $S$ is a commutative subsemigroup and hence a meet semilattice, with the partial order given by $e\leq f$ if and only if $ef=e$, and the meet by $e\wedge f=ef$. The order extends to $S$ by defining $s\leq t$ if $s=te$ for some $e\in E(S)$. If $e\in E(S)$, then $G_{e}=\{s\in S\mid s^{*}s=e=ss^{*}\}$ is a group with identity $e$ known as the maximal subgroup of $S$ at $e$.
A topological groupoid is a groupoid $\mathscr{G}$ endowed with a locally compact topology such that composition and inversion are continuous. We do not assume that $\mathscr{G}$ is Hausdorff in general, but we do assume that the unit space $\mathscr{G}^{(0)}=\{\gamma\gamma^{-1}\mid\gamma\in\mathscr{G}\}$ is Hausdorff with the relative topology inherited from $\mathscr{G}$. Since we do not assume that $\mathscr{G}$ is Hausdorff, locally compact is defined as every point in $\mathscr{G}$ having a compact neighborhood, where compact means Hausdorff and open covers have finite subcovers. Following the Bourbaki tradition, we say that a space $X$ is quasi-compact if every open covering of $X$ has a finite subcovering (but with no assumption of being Hausdorff). Note that a finite union of quasi-compact sets is quasi-compact and a closed subset of a quasi-compact set is quasi-compact.
An étale groupoid is a topological groupoid $\mathscr{G}$ such that the range map $\mathop{\boldsymbol{r}}\colon\mathscr{G}\to\mathscr{G}^{(0)}$ (or equivalently, the source map $\mathop{\boldsymbol{s}}$) is a local homeomorphism. A bisection of an étale groupoid $\mathscr{G}$ is an open subset $U\subseteq\mathscr{G}$ such that the restrictions $\mathop{\boldsymbol{r}}|_{U}$ and $\mathop{\boldsymbol{s}}|_{U}$ are homeomorphisms onto an open set in $\mathscr{G}^{(0)}$. In particular, bisections are always Hausdorff subspaces. An étale groupoid $\mathscr{G}$ is an ample groupoid if $\mathscr{G}^{(0)}$ is a locally compact Hausdorff space with a basis of compact open sets.
@@
$\mathscr{G}^{x}_{x}=\{\gamma\in\mathscr{G}\mid\mathop{\boldsymbol{s}}(\gamma)=x=\mathop{\boldsymbol{r}}(\gamma)\}.$
@@
If $X\subseteq\mathscr{G}^{(0)}$, then $\mathscr{G}|_{X}$ is the subgroupoid of all arrows $\gamma$ with $\mathop{\boldsymbol{s}}(\gamma),\mathop{\boldsymbol{r}}(\gamma)\in X$, i.e., $\mathscr{G}|_{X}=\mathop{\boldsymbol{s}}^{-1}(X)\cap\mathop{\boldsymbol{r}}^{-1}(X)$. In particular, if $X$ is open (respectively, closed) in $\mathscr{G}^{(0)}$, then $\mathscr{G}|_{X}$ is open (respectively, closed) in $\mathscr{G}$.
The orbit of $x\in\mathscr{G}^{(0)}$ is the set $\mathcal{O}_{x}$ consisting of all $y\in\mathscr{G}^{(0)}$ such that there is an arrow $\gamma\colon x\to y$. Note that $\mathcal{O}_{x}=\mathop{\boldsymbol{r}}(\mathop{\boldsymbol{s}}^{-1}(x))=\mathop{\boldsymbol{s}}(\mathop{\boldsymbol{r}}^{-1}(x))$. The set of all orbits of $\mathscr{G}$ is denoted $\mathscr{G}^{(0)}/\mathscr{G}$. A subset $X$ of $\mathscr{G}^{(0)}$ is invariant if it is a union of orbits.
@@
respectively. The semilattice of idempotents of $\Gamma_{c}(\mathscr{G})$ is given by the set of compact open subsets of $\mathscr{G}^{(0)}$.
2.2. Algebras of ample groupoids
Fix an ample groupoid $\mathscr{G}$ and a field $K$. The Steinberg algebra of $\mathscr{G}$ over $K$ is
$K\mathscr{G}=\mathrm{span}_{K}\{1_{U}\mid U\in\Gamma_{c}(\mathscr{G})\},$
@@ 2.2. Algebras of ample groupoids
$f\ast g(\gamma)=\sum_{\mathop{\boldsymbol{s}}(\gamma)=\mathop{\boldsymbol{s}}\alpha)}f(\gamma\alpha^{-1})g(\alpha)=\sum_{\alpha\beta=\gamma}f(\alpha)g(\beta).$
Notice that $1_{U}*1_{V}=1_{UV}$ for all $U,V\in\Gamma_{c}(\mathscr{G})$. Thus $K\mathscr{G}$ is a quotient of the semigroup algebra $K\Gamma_{c}(\mathscr{G})$. If $K$ is a field with involution $\sigma$ (possibly trivial), then $K\mathscr{G}$ is a $\ast$-algebra with involution given by $f^{\ast}(\gamma)=\sigma(f(\gamma^{-1}))$. We shall only use this in the case $K=\mathbb{C}$ and $\sigma$ is complex conjugation. Note that $K\mathscr{G}$ is unital if and only if $\mathscr{G}^{(0)}$ is compact. See [31] for details.
If $U$ is an open invariant subset of $\mathscr{G}^{(0)}$ and $X=\mathscr{G}^{(0)}\setminus U$ is the complementary closed invariant subspace, then $I=K\mathscr{G}|_{U}$ is an ideal of $K\mathscr{G}$ and $K\mathscr{G}/I\cong K\mathscr{G}|_{X}$ via restriction of functions by [10, Pages 1596–1597], a fact we shall use in the sequel without comment.
@@ 2.3. Groupoid of germs and the universal groupoid of an inverse semigroup
Fix an inverse semigroup $S$ and let $E$ denote the semilattice of idempotents of $S$ (since $S$ will be fixed throughout this subsection). Let $X$ be a locally compact totally disconnected space. We denote by $I_{X}$ the inverse semigroup of all partial homeomorphisms of $X$ with compact open domain and range. An action of $S$ on $X$ is a homomorphism $\varphi\colon S\to I_{X}$ such that if $X_{e}$ denotes the domain of an idempotent $e$, then $\bigcup_{e\in E}X_{e}=X$; this last condition says that the action is non-degenerate. Notice that $\psi_{s}\colon X_{s^{*}s}\to X_{ss^{*}}$ for each $s\in S$. We put $sx=\psi_{s}(x)$ for the action when no confusion can arise.
@@
The unit space $\mathscr{G}_{S}^{(0)}=\{[e,x]\mid x\in X_{e}\}$ is identified with $X$ via the homeomorphism $[e,x]\mapsto x$. From this point of view, the domain and range maps are given by $\mathop{\boldsymbol{s}}([s,x])=x$ and $\mathop{\boldsymbol{r}}([s,x])=sx$.
@@
is a homeomorphism, and the map $s\mapsto\alpha_{s}$ defines an action $\alpha\colon S\to I_{\widehat{E}}$ of $S$ on $\widehat{E}$ (see [29, Proposition 4.3.2] [^1] Note that Paterson works with right actions in [29].).
The universal groupoid $\mathscr{G}_{S}$ associated with $S$ is the groupoid of germs $\mathscr{G}_{S}$ of the action of $S$ on $\widehat{E}$. For $s\in S$, define $\psi(s)=(s,D_{s^{*}s})$. Then $\psi$ is a homomorphism of $S$ into $\Gamma_{c}(\mathscr{G}_{S})$[29, Theorem 3.3.2] [^2] In fact, $\psi$ is a restriction of Paterson’s $\psi_{X}$ in [29, Theorem 3.3.2]. In fact, $\psi$ is injective, and extends to an isomorphism $\widehat{\psi}\colon KS\to K\mathscr{G}_{S}$; see [31, Theorem 6.3].

### v3

@@ Abstract.
We completely characterize when the algebra of an ample groupoid with coefficients in an arbitrary unital ring is von Neumann regular and, more generally, when the algebra of a graded ample groupoid is graded von Neumann regular. Our main application is to resolve the question, open since 1970, of when the algebra of an inverse semigroup is von Neumann regular. As applications we recover known results on regularity and graded regularity of Leavitt path algebras, and prove a number of new results, in particular concerning graded regularity of algebras of Deaconu-Renault groupoids and Nekrashevych-Exel-Pardo algebras of self-similar groups.
@@ 1. Introduction
A famous result of Connell [12] says that a group ring $RG$, over a unital ring $R$, is regular if and only if $R$ is regular, $G$ is locally finite and the order of every finite subgroup of $G$ is invertible in $R$. In particular, this says that regularity is a rare feature for groups rings. Leavitt path algebras [1] are also rarely regular: they are regular if and only if the (directed) graph defining the Leavitt path algebra is acyclic [3]. However, every Leavitt path algebra over a field is graded von Neumann regular by a result of Hazrat [20]. A graded ring $R$ is graded regular if the semigroup of homogeneous elements is regular. This has an important impact on the graded $K$-theory of Leavitt path algebras [21].
Okniński tried to generalize Connell’s results to inverse semigroup algebras (see his book [29]). He was able to prove that if $K$ is a field of characteristic 00 and $S$ is an inverse semigroup, then $KS$ is regular if and only if $S$ is locally finite. In positive characteristic he was able to give necessary conditions: $S$ is periodic, each maximal subgroup of $S$ is locally finite and the characteristic of $K$ cannot divide the order of any element of a maximal subgroup. It had been proven earlier by Weissglass [34] that if $R$ is regular, $S$ is locally finite and the order of every finite subgroup of $S$ is invertible in $R$, then $RS$ is regular. But being periodic with locally finite maximal subgroups is a strictly weaker condition than being locally finite. The reason that Okniński was able to obtain a complete characterization in characteristic 00 is that he could embed $\mathbb{C}S$ into the Banach algebra $\ell_{1}(S)$ and use some analysis.
@@
Our main result is a complete characterization of regularity of $R\mathscr{G}$ over any base unital ring $R$. Following the Bourbaki tradition, we say a space is quasi-compact if every open covering has a finite subcover and reserve “compact” to mean quasi-compact and Hausdorff.
Theorem A.
Let $\mathscr{G}$ be an ample groupoid and $R$ a unital ring. Then $R\mathscr{G}$ is von Neumann regular if and only if:
  1. (1)
    $R$ is von Neumann regular;
  2. (2)
    $\mathscr{G}$ is a directed union of quasi-compact open subgroupoids;
  3. (3)
    the order of each finite subgroup of an isotropy group of $\mathscr{G}$ is invertible in $R$.
For principal, second countable Hausdorff groupoids the second condition is equivalent to approximate finiteness in the sense of [18], hence we call such groupoids approximately quasi-compact.
As a consequence, we resolve the problem of characterizing regular inverse semigroup algebras, dating back to the 1970 paper of Weissglass [34].
Theorem B.
Let $S$ be an inverse semigroup and $R$ a unital ring. Then $RS$ is von Neumann regular if and only if:
  1. (1)
    $R$ is von Neumann regular;
  2. (2)
    $S$ is locally finite;
  3. (3)
    the order of each finite subgroup of $S$ is invertible in $R$.
The sufficiency of these conditions goes back to Weissglass [34].
We also consider graded regularity of groupoid algebras for ample groupoids equipped with a locally constant cocycle to a group. We prove the algebra is graded regular if and only if the homogeneous component of the identity is regular. Since the homogeneous component of the identity is the algebra of the clopen subgroupoid of elements mapping to the identity under the cocycle, our results for regularity of groupoid algebras apply. From this we easily recover Hazrat’s result on graded regularity of Leavitt path algebras [20]. We extend these results to algebras of Deaconu-Renault groupoids under mild hypotheses. These include algebras of row finite higher rank graphs. We consider further graded regularity of partial group skew products, Leavitt path algebras of labeled graphs and Nekrashevych-Exel-Pardo algebras of self-similar graphs.
The paper is organized as follows. Section 2 is devoted to preliminaries concerning inverse semigroups, ample groupoids and regular rings. Then Section 3 proves our main results on regularity of ample groupoid algebras. Section 4 addresses the question of graded regularity of the algebra of an ample groupoid equipped with a locally constant cocycle. The final section, applies the main results to study regularity and graded regularity for the algebras of a number of well-studied families of ample groupoids. We added an appendix giving a groupoid-free proof of Theorem B for specialists in semigroup theory, although the proof is merely a retranslation of the groupoid proof.
Acknowledgments
We thank Gilles de Castro for showing Example 5.11 to us.
@@ 2.1. Inverse semigroups and groupoids
The reader is referred to [23] for details on the theory of inverse semigroups. A semigroup is a nonempty set together with an associative binary operation. An inverse semigroup is a semigroup $S$ such that for every $s\in S$ there is a unique $s^{*}\in S$ such that $s^{*}ss^{*}=s^{*}$ and $ss^{*}s=s$. Notice that $ss^{*},s^{*}s$ are idempotents for all $s\in S$. The set $E(S)$ of idempotents of $S$ is a commutative subsemigroup and hence a meet semilattice with the partial order given by $e\leq f$ if and only if $ef=e$, and the meet by $e\wedge f=ef$. The order extends to $S$ by defining $s\leq t$ if $s=te$ for some $e\in E(S)$. If $e\in E(S)$, then $G_{e}=\{s\in S\mid s^{*}s=e=ss^{*}\}$ is a group with identity $e$ known as the maximal subgroup of $S$ at $e$.
A topological groupoid is a groupoid $\mathscr{G}$ endowed with a locally compact topology such that composition and inversion are continuous. All groupoids are assumed nonempty. We do not assume that $\mathscr{G}$ is Hausdorff in general, but we do assume that the unit space $\mathscr{G}^{(0)}=\{\gamma\gamma^{-1}\mid\gamma\in\mathscr{G}\}$ is Hausdorff with the relative topology inherited from $\mathscr{G}$. Since we do not assume that $\mathscr{G}$ is Hausdorff, locally compact is defined as every point in $\mathscr{G}$ having a compact neighborhood, where compact means Hausdorff and open covers have finite subcovers. Following the Bourbaki tradition, we say that a space $X$ is quasi-compact if every open covering of $X$ has a finite subcovering (but with no assumption of being Hausdorff). Note that a finite union of quasi-compact sets is quasi-compact and a closed subset of a quasi-compact set is quasi-compact.
An étale groupoid is a topological groupoid $\mathscr{G}$ such that the range map $\mathop{\boldsymbol{r}}\nolimits\colon\mathscr{G}\to\mathscr{G}^{(0)}$ (or equivalently, the source map $\mathop{\boldsymbol{s}}$) is a local homeomorphism. A bisection of an étale groupoid $\mathscr{G}$ is an open subset $U\subseteq\mathscr{G}$ such that the restrictions $\mathop{\boldsymbol{r}}\nolimits|_{U}$ and $\mathop{\boldsymbol{s}}|_{U}$ are homeomorphisms onto an open set in $\mathscr{G}^{(0)}$. In particular, bisections are always Hausdorff subspaces. An étale groupoid $\mathscr{G}$ is an ample groupoid if $\mathscr{G}^{(0)}$ is a locally compact Hausdorff space with a basis of compact open sets.
@@
$\mathscr{G}^{x}_{x}=\{\gamma\in\mathscr{G}\mid\mathop{\boldsymbol{s}}(\gamma)=x=\mathop{\boldsymbol{r}}\nolimits(\gamma)\}.$
@@
If $X\subseteq\mathscr{G}^{(0)}$, then $\mathscr{G}|_{X}$ is the subgroupoid of all arrows $\gamma$ with $\mathop{\boldsymbol{s}}(\gamma),\mathop{\boldsymbol{r}}\nolimits(\gamma)\in X$, i.e., $\mathscr{G}|_{X}=\mathop{\boldsymbol{s}}^{-1}(X)\cap\mathop{\boldsymbol{r}}\nolimits^{-1}(X)$. In particular, if $X$ is open (respectively, closed) in $\mathscr{G}^{(0)}$, then $\mathscr{G}|_{X}$ is open (respectively, closed) in $\mathscr{G}$.
The orbit of $x\in\mathscr{G}^{(0)}$ is the set $\mathcal{O}_{x}$ consisting of all $y\in\mathscr{G}^{(0)}$ such that there is an arrow $\gamma\colon x\to y$. Note that $\mathcal{O}_{x}=\mathop{\boldsymbol{r}}\nolimits(\mathop{\boldsymbol{s}}^{-1}(x))=\mathop{\boldsymbol{s}}(\mathop{\boldsymbol{r}}\nolimits^{-1}(x))$. The set of all orbits of $\mathscr{G}$ is denoted $\mathscr{G}^{(0)}/\mathscr{G}$. A subset $X$ of $\mathscr{G}^{(0)}$ is invariant if it is a union of orbits.
@@
respectively. The semilattice of idempotents of $\Gamma_{c}(\mathscr{G})$ is given by the set of compact open subsets of $\mathscr{G}^{(0)}$. Notice that if $\mathscr{G}$ is an ample group bundle, then $\Gamma_{c}(\mathscr{G})$ is a Clifford semigroup, as $\mathop{\boldsymbol{s}}(U)=\mathop{\boldsymbol{r}}\nolimits(U)$ for all $U\in\Gamma_{c}(\mathscr{G})$.
2.2. Algebras of inverse semigroups and ample groupoids
Fix a unital ring $R$ (not necessarily commutative). If $S$ is an inverse semigroup, then the inverse semigroup ring $RS$ is the free left $R$-module with basis $S$ and with multiplication
$\sum_{s\in S}c_{s}s\cdot\sum_{t\in S}d_{t}t=\sum_{s,t\in S}c_{s}d_{t}st.$
Fix an ample groupoid $\mathscr{G}$. Then the Steinberg algebra of $\mathscr{G}$ over $R$ is the left $R$-module
$R\mathscr{G}=\mathrm{span}_{R}\{1_{U}\mid U\in\Gamma_{c}(\mathscr{G})\}\subseteq R^{\mathscr{G}},$
@@ 2.2. Algebras of inverse semigroups and ample groupoids
$f\ast g(\gamma)=\sum_{\mathop{\boldsymbol{s}}(\gamma)=\mathop{\boldsymbol{s}}(\alpha)}f(\gamma\alpha^{-1})g(\alpha)=\sum_{\alpha\beta=\gamma}f(\alpha)g(\beta).$
Notice that $1_{U}*1_{V}=1_{UV}$ for all $U,V\in\Gamma_{c}(\mathscr{G})$. Thus $R\mathscr{G}$ is a quotient of the semigroup algebra $R\Gamma_{c}(\mathscr{G})$. Note that $R\mathscr{G}$ is unital if and only if $\mathscr{G}^{(0)}$ is compact. See [33] for details (where the assumption that $R$ is commutative is not necessary).
If $U$ is an open invariant subset of $\mathscr{G}^{(0)}$ and $X=\mathscr{G}^{(0)}\setminus U$ is the complementary closed invariant subspace, then $I=R\mathscr{G}|_{U}$ is an ideal of $R\mathscr{G}$ and $R\mathscr{G}/I\cong R\mathscr{G}|_{X}$ via restriction of functions by [10, Pages 1596–1597] (note an omission in the argument is resolved in [24, Corollar 4.4]), a fact we shall use in the sequel without comment.
@@ 2.3. Groupoid of germs and the universal groupoid of an inverse semigroup
Fix an inverse semigroup $S$ and let $E$ denote the semilattice of idempotents of $S$ (since $S$ will be fixed throughout this subsection). Let $X$ be a locally compact totally disconnected space. We denote by $I_{X}$ the inverse semigroup of all partial homeomorphisms of $X$ with compact open domain and range. An action of $S$ on $X$ is a homomorphism $\varphi\colon S\to I_{X}$ such that if $X_{e}$ denotes the domain of an idempotent $e$, then $\bigcup_{e\in E}X_{e}=X$; this last condition says that the action is non-degenerate. Notice that $\varphi_{s}\colon X_{s^{*}s}\to X_{ss^{*}}$ for each $s\in S$. We put $sx=\varphi_{s}(x)$ for the action when no confusion can arise.
@@
The unit space $\mathscr{G}_{S}^{(0)}=\{[e,x]\mid x\in X_{e}\}$ is identified with $X$ via the homeomorphism $[e,x]\mapsto x$. From this point of view, the domain and range maps are given by $\mathop{\boldsymbol{s}}([s,x])=x$ and $\mathop{\boldsymbol{r}}\nolimits([s,x])=sx$.
@@
is a homeomorphism, and the map $s\mapsto\alpha_{s}$ defines an action $\alpha\colon S\to I_{\widehat{E}}$ of $S$ on $\widehat{E}$ (see [30, Proposition 4.3.2] [^1] Note that Paterson works with right actions in [30].).
The universal groupoid $\mathscr{G}_{S}$ associated with $S$ is the groupoid of germs $\mathscr{G}_{S}$ of the action of $S$ on $\widehat{E}$. For $s\in S$, define $\psi(s)=(s,D_{s^{*}s})$. Then $\psi$ is a homomorphism of $S$ into $\Gamma_{c}(\mathscr{G}_{S})$[30, Theorem 3.3.2] [^2] In fact, $\psi$ is a restriction of Paterson’s $\psi_{X}$ in [30, Theorem 3.3.2]. In fact, $\psi$ is injective, and extends to an isomorphism $\widehat{\psi}\colon RS\to R\mathscr{G}_{S}$; see [33, Theorem 6.3] where the assumption that $R$ is commutative is unnecessary.